# A notion of fractality for a class of states and noncommutative relative distance zeta functional

### v3

@@ Abstract
In this work, we first recall the definition of the relative distance zeta function in [42, 43, 44, 46, 47] and slightly generalize this notion from sets to probability measures, and then move on to propose a novel definition a relative distance (and tube) zeta functional for a class of states over a C* algebra. With such an extension, we look into the chance to define relative Minkowski dimensions in this context, and explore the notion of fractality for this class of states. Relative complex dimensions as poles of this newly proposed relative distance zeta functional, as well as its geometric and transformation properties, decomposition rules and properties that respects tensor products are discussed. We then explore some examples that possess fractal properties with this new zeta functional and propose functional equations similar to [11].
@@ 1 Introduction
In this work, we make a pioneering step to explore the possibility of extending the definition of the relative distance (and tube) zeta function in [25, 26, 27, 41, 42, 43, 44, 45, 46, 47, 48, 49] to the noncommutative setting, more specifically to a class of states over a C* algebra. (A relevent but slightly different generalization of fractal zeta functions onto (Ahlfors) metric measure spaces are also explored in [8, 52], see also [43].) We actualize this by first slightly extending this notion from sets to probability measures in the Euclidean space as in (1), and view them as a special case of states over a commutative C* algebra. Then we observe a consistent generalization can be made to provide the definition of a relative distance (and tube) zeta functional as in (5) for a special class of states. Along this line, we may also define tubular neighborhood measures and functionals which help us define relative Minkowski dimensions in this context. Relative complex dimensions can then be defined as poles of this newly proposed relative distance zeta functional, and can be regarded as a generaliztion of dimensions in view of the fact that under mild hypotheses the relative upper Minkowski dimension is always contained in the set of relative complex dimensions. Similar to the commutative case, the presence of the imaginary parts in the relative complex dimensions, i.e. the poles of the relative distance zeta functional, captures the periodicity in the logarithmic scale, i.e. the multiplicative periodicity, of the tubular neighborhood functionals. This therefore gives us the right tools to explore the notion of fractality for our class of states over a C* algebra, which, being a limit of convex combinations of pure states via Choquet’s theorem, may be regarded as a noncommutative generalization of geometric objects when we recall the fact that pure states over commutative C* algebra are always Dirac measures over points in a space. Additional geometric and transformation properties, decomposition rules and properties that respects tensor products are explored such as functional equations similar to those introduced in [11] could be obtained, which could possibly offer us a better description of the relative complex dimensions of the states considered. Functional equations satisfied by the geometric, fractal and distance zeta function with rescaling as in [11] were previously also introduced and used earlier in the following references, including [12, 13, 14, 15, 16, 21, 25, 26, 27, 34, 35, 36, 37, 42, 43, 45]
@@
As an important remark to further understand the motivtion of this work and its physical relevance, we would like to turn our attention to, and compare with, another major direction of fractal studies in the noncommutative setup, notably [4, 17] in which the authors perform a quantization of fractals via the study of their spectral triple through the lens of Gromov-Hausdorff propinquity over quantum metric space [53, 54]. Other works that aims for a quantization of fractals via the study of their spectral triple include [5, 18, 57, 60]. The difference between our work and thes aforementioned approaches is majorly on the respective motivtions: their works were to look for a noncommutative characterization of (known) fractals (via a spectral triple, for instance), whereas our work and approach is to explore the other way around, asking the question of what might be considered as a fractal in a noncommutative setting, more concretely in the context of the C*-algebraic formulation of quantum mechanics where physical observables are their self-adjoint elements. To understand more concretely what we mean here, let us recall that, in this setup, states of a C*-algebra can be regarded as a generalization of the notion of density matrices in quantum mechanics (which represent mixed “quantum states” in the physical system in quantum entanglement or statistical ensemble) and they now correspond to physical states as a mapping from physical observables (as self-adjoint elements) to their expected measurement outcomes (as real or complex numbers.) Therefore asking for a characterization of fractality of states is equivalent to asking for such a characterization of a quantum system given only expected measurement outcomes of the relevant restricted set of physical observables (viewed as a geometric object). A growing net of tubular neighborhood functionals then represents having increasingly more information of expected measurement outcomes over an increasingly larger set of physical observables which contains our relevant restricted set to be studied, the fractality of which is now described by the periodicity in the logarithmic scale/multiplicative periodicity to be observed when the set of physical observables from where information for expected measurement outcomes can be obtained grows.
@@ 2.1 The commutative relative distance zeta function
where $d(x,y)=\sqrt{\sum_{i=1}^{d}(x_{i}-y_{i})^{2}}$ is the Euclidean distance and $\text{spt}(\mu)$ denotes the support of $\mu$:
$\text{spt}(\mu):=\overline{\{A\in\mathcal{B}\,:\,\mu(A)\neq 0\ \}}$
where $\mathcal{B}$ is the Borel sigma algebra. Then we may consider, as in [42, 43, 44, 46, 47], for another given Borel probability measure $\nu\in\mathcal{M}(\mathbb{R}^{d})$, the relative distance zeta function.
@@
Furthermore, if the relative box counting dimension of $\mu$ with respect to $\nu$, i.e. $\text{dim}_{B}(\mu,\nu)$, exists and $\text{dim}_{B}(\mu,\nu)<d$, and moreover if $\mathcal{M}^{\text{dim}_{B}(\mu,\nu)}_{*}(\mu,\nu)>0$, then $\zeta_{\mu}(s,\nu)\rightarrow\infty$ when $s\rightarrow\text{dim}_{B}(\mu,\nu)^{+}$ for $s\in\mathbb{R}$.
@@
The first part and third part of the lemma come from an equivalence of norms in the finite dimension, according to which
@@ 2.2 The noncommutative relative distance zeta functional
where $\mu_{\tau}\in\mathcal{M}(P(H))$ is a Borel probability measure on $P(H):=\partial B_{1}^{H}(0)/\{\pm 1\}$; and we denote $\langle w,X\,w\rangle_{H}:=\langle r,X\,r\rangle_{\tilde{H}}$ whenever $r\in w$. When $n=\text{dim}_{\mathbb{R}}(H)$ can be possibly $\infty$, we may on the other hand consider the GNS representation of $\mathcal{L}(H)$ with $\tilde{H}:=\oplus_{l\in\mathcal{PS}(\mathcal{L}(H))}\,\mathcal{L}(H)/(l(a^{*}a)=0)$ (which is not separable) where we now denote $\mathcal{PS}(\mathcal{L}(H))$ as the set of pure states of $\mathcal{L}(H)$. With that, we obtain the following representation of $\tau$ via the Choquet theorem:
@@
Hence, when $n=\text{dim}_{\mathbb{R}}(H)=2\,\text{dim}_{\mathbb{C}}(H)<\infty$, there is a canonical map between the coordinate vectors of P(H~)P($\tilde${H}) and the elements on $P(H)$, and the construction of $\tilde{H}$ is unnecessary (especially noticing that $\tilde{H}$ is way bigger). However, when $n=\text{dim}_{\mathbb{R}}(H)=\infty$, we cannot construct this map, as it is well-known that there are states $l$ such that $l(\mathcal{K}(H))=0$, where $\mathcal{K}(H)$ is the space of compact operators. (Meanwhile, as a possibly slightly confusing remark, if we start with an arbitrary $C^{*}$-algebra, we may take $H$ itself as the GNS representation of the $C^{*}$-algebra right from the start, and a state over that $C^{*}$-algebra can already be represented as an integral previously described (noting $n=\text{dim}_{\mathbb{R}}(H)=2\,\text{dim}_{\mathbb{C}}(H)=\infty$) i.e.
$\tau(X)=\int_{P(H)}\langle w,X\,w\rangle_{H}\,\mu_{\tau}(dw)$
@@
Motivated by the previous discussion as well as the commutative case, let us define the following:
@@
We notice that this definition is not restricted to the case when $n:=\text{dim}_{\mathbb{R}}(H)<\infty$, and the relative distance zeta functional can be explicitly written as
@@
Therefore our definition of (5) is not (interpreted in any sense as) a distance zeta function treating $\mathcal{L}(H)$ as metric space. In fact, when $n$ is finite, our choice of having an infrimum over $\sigma(\mathcal{E}(X-Y))$ instead of a supremum compensates for the dimensionality effect within one variable $X_{i}$ from the aforementioned dimensionality argument. We postulate that we always have $\overline{\text{dim}}_{B}(\tau,\nu)$ on the strip $\{0\leq\Re(s)\leq d\}$, irrespective of whether $n$ is finite or infinite.
@@
Furthermore, if the relative box counting dimension of $\tau$ with respect to $\nu$, i.e. $\text{dim}_{B}(\tau,\nu)$, exists and $\text{dim}_{B}(\tau,\nu)<d$, and moreover if $\mathcal{M}^{\text{dim}_{B}(\tau,\nu)}_{*}(\tau,\nu)>0$, then $\|\zeta_{\tau}(s,\nu)\|\rightarrow\infty$ when $s\rightarrow\text{dim}_{B}(\tau,\nu)^{+}$ for $s\in\mathbb{R}$.
@@
Given an arbitrary $\nu\in\tilde{\mathcal{M}}(\overline{\mathcal{A}\langle X\rangle})$, and for a fixed $\delta>0$, the $\delta$-tubular-neighborhood linear functional of $\tau\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ with respect to $\nu$ always satisfies $\text{spt}(\mu_{\nu^{\tau,\delta}})\subset\{(X,\xi)\in[\mathfrak{sa}\,\mathcal{L}(H)]^{d}\times P(\tilde{H}):d(X,\tau)\leq t\}$, and therefore $\zeta_{\tau}(s,\nu^{\tau,\delta})$ is well-defined and holomorphic whenever $\Re(s)>\overline{\text{dim}}_{B}(\mu,\nu^{\tau,\delta})$. Also, the above representation of $\zeta_{\tau}(s,\nu^{\tau,\delta})$ in the lemma shows that, given $g\in\overline{\mathcal{A}\langle X\rangle}$
@@
A quick remark is now in force, that it is now unclear if the relative complex dimension $\mathcal{D}_{\mathbb{C}}(\tau,\nu)$, as a union of $\mathcal{D}_{\mathbb{C}}(\tau,\nu)[g]$ over all $g\in\overline{\mathcal{A}\langle X\rangle}$, is in general a discrete set.
@@ 3.2 Decomposition rules
    Suppose $\tau\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ can be decomposited into $\tau=a_{1}\tau_{1}+a_{2}\tau_{2}$ with $\tau_{1},\tau_{2}\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ and $a_{1}+a_{2}=1$. Then, for $\nu\in\tilde{\mathcal{M}}(\overline{\mathcal{A}\langle X\rangle})$, then whenever $\overline{\text{dim}}_{B}(\tau_{2},\nu)<\overline{\text{dim}}_{B}(\tau,\nu)$, and for all $s\in\Omega_{\left[\zeta_{\tau_{1}}(\,\cdot\,,\nu)\right](g)}\bigcap\{s\in\mathbb{C}\,:\,\Re(s)>\overline{\text{dim}}_{B}(\tau_{2},\nu)\}$,
@@
    Suppose $\nu\in\tilde{\mathcal{M}}(\overline{\mathcal{A}\langle X\rangle})$ is given, and $\tau\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ can be decomposited into $\tau=a_{1}\tau_{1}+a_{2}\tau_{2}$ with $\tau_{1},\tau_{2}\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ and $a_{1}+a_{2}=1$ such that for some $\epsilon>0$,
@@
where we know apriori that the right hand side of the second inequality is finite because $\left\|\nu\right\|<\infty$. Therefore we have
@@ 3.3 Zeta functional over tensor states and decomposition rules
$\nu(f)=\int_{[\mathfrak{sa}\,\mathcal{L}(H)]^{\sum_{i=1}^{k}d_{i}}\times[P(\tilde{H})]^{k}}\,\left\langle f(X_{1},...,X_{k})\,\otimes_{i=1}^{k}\xi_{i}\,\,,\,\otimes_{i=1}^{k}\xi_{i}\right\rangle_{\tilde{H}^{\otimes^{k}}}\,\mu_{\nu}(d\xi_{1},...d\xi_{k},dX_{1},...,dX_{k})$
@@
Furthermore, if the relative box counting dimension of $\tau$ with respect to $\nu$, i.e. $\text{dim}_{B}(\tau,\nu)$, exists and $\text{dim}_{B}(\tau,\nu)<d$, and moreover if $\mathcal{M}^{\text{dim}_{B}(\tau,\nu)}_{*}(\tau,\nu)>0$, then $\|\zeta_{\tau}(s,\nu)\|\rightarrow\infty$ when $s\rightarrow\text{dim}_{B}(\tau,\nu)^{+}$ for $s\in\mathbb{R}$.
@@
$\quad\quad\quad\quad\quad\prod_{i=1}^{k}\langle\xi_{i},g_{i}(X_{i})\,\xi_{i}\rangle_{\tilde{H}}\,(\otimes_{i=1}^{k}\mu_{\nu_{i}})(d\xi_{1},...d\xi_{k},dX_{1},...,dX_{k})\,.$
@@
The first part and third part of the lemma comes from an equivalence of norm in the finite dimensions

### v4

@@ Abstract
In this work, we first recall the definition of the relative distance zeta function in [42, 43, 44, 46, 47] and slightly generalize this notion from sets to probability measures, and then move on to propose a novel definition a relative distance (and tube) zeta functional for a class of states over a C* algebra. With such an extension, we look into the chance to define relative Minkowski dimensions in this context, and explore the notion of fractality for this class of states. Relative complex dimensions as poles of this newly proposed relative distance zeta functional, as well as its geometric and transformation properties, decomposition rules and properties that respects tensor products are discussed. We then explore some examples that possess fractal properties with this new zeta functional and propose functional equations similar to [11, 35, 36, 42].
@@ 1 Introduction
In this work, we make a pioneering step to explore the possibility of extending the definition of the relative distance (and tube) zeta function in [25, 26, 27, 41, 42, 43, 44, 45, 46, 47, 48, 49] to the noncommutative setting, more specifically to a class of states over a C* algebra. (A relevent but slightly different generalization of fractal zeta functions to (Ahlfors) metric measure spaces are also explored in [8, 52], see also [43].) We actualize this by first slightly extending this notion from sets to probability measures in the Euclidean space as in (1), and view them as a special case of states over a commutative C* algebra. Then we observe a consistent generalization can be made to provide the definition of a relative distance (and tube) zeta functional as in (5) for a special class of states. Along this line, we may also define tubular neighborhood measures and functionals which help us define relative Minkowski dimensions in this context. Relative complex dimensions can then be defined as poles of this newly proposed relative distance zeta functional, and can be regarded as a generaliztion of dimensions in view of the fact that under mild hypotheses the relative upper Minkowski dimension is always contained in the set of relative complex dimensions. Similar to the commutative case, the presence of the imaginary parts in the relative complex dimensions, i.e. the poles of the relative distance zeta functional, captures the periodicity in the logarithmic scale, i.e. the multiplicative periodicity, of the tubular neighborhood functionals. This therefore gives us the right tools to explore the notion of fractality for our class of states over a C* algebra, which, being a limit of convex combinations of pure states via Choquet’s theorem, may be regarded as a noncommutative generalization of geometric objects when we recall the fact that pure states over commutative C* algebra are always Dirac measures over points in a space. Additional geometric and transformation properties, decomposition rules and properties that respects tensor products are explored such as functional equations similar to those introduced in [11] could be obtained, which could possibly offer us a better description of the relative complex dimensions of the states considered. Functional equations satisfied by the geometric, fractal and distance zeta function with rescaling as in [11] were previously also introduced and used earlier in the following references, including [12, 13, 14, 15, 16, 21, 25, 26, 27, 34, 35, 36, 37, 42, 43, 45].
@@
As an important remark to further understand the motivation of this work and its physical relevance, we would like to turn our attention to, and compare with, another major direction of fractal studies in the noncommutative setup, notably [4, 17] in which the authors perform a quantization of fractals via the study of their spectral triple through the lens of Gromov-Hausdorff propinquity over quantum metric space [53, 54]. Other works that aims for a quantization of fractals via the study of their spectral triple include [5, 18, 57, 60]. The difference between our work and thes aforementioned approaches is majorly on the respective motivations: their works were to look for a noncommutative characterization of (known) fractals (via a spectral triple, for instance), whereas our work and approach is to explore the other way around, asking the question of what might be considered as a fractal in a noncommutative setting, more concretely in the context of the C*-algebraic formulation of quantum mechanics where physical observables are their self-adjoint elements. To understand more concretely what we mean here, let us recall that, in this setup, states of a C*-algebra can be regarded as a generalization of the notion of density matrices in quantum mechanics (which represent mixed “quantum states” in the physical system in quantum entanglement or statistical ensemble) and they now correspond to physical states as a mapping from physical observables (as self-adjoint elements) to their expected measurement outcomes (as real or complex numbers.) Therefore asking for a characterization of fractality of states is equivalent to asking for such a characterization of a quantum system given only expected measurement outcomes of the relevant restricted set of physical observables (viewed as a geometric object). A growing net of tubular neighborhood functionals then represents having increasingly more information of expected measurement outcomes over an increasingly larger set of physical observables which contains our relevant restricted set to be studied, the fractality of which is now described by the periodicity in the logarithmic scale/multiplicative periodicity to be observed when the set of physical observables from where information for expected measurement outcomes can be obtained grows.
@@ 2.1 The commutative relative distance zeta function
where $d(x,y)=\sqrt{\sum_{i=1}^{d}(x_{i}-y_{i})^{2}}$ is the Euclidean distance and $\text{spt}(\mu)$ denotes the support of $\mu$. Then we may consider, as in [42, 43, 44, 46, 47], for another given Borel probability measure $\nu\in\mathcal{M}(\mathbb{R}^{d})$, the relative distance zeta function.
@@
Furthermore, if the relative box counting dimension of $\mu$ with respect to $\nu$, i.e. $\text{dim}_{B}(\mu,\nu)$, exists and $\text{dim}_{B}(\mu,\nu)<d$, and moreover, if $\mathcal{M}^{\text{dim}_{B}(\mu,\nu)}_{*}(\mu,\nu)>0$, then $\zeta_{\mu}(s,\nu)\rightarrow\infty$ when $s\rightarrow\text{dim}_{B}(\mu,\nu)^{+}$ for $s\in\mathbb{R}$.
@@
The first part and third part of the lemma come from an equivalence of norms in finite dimension, according to which
@@ 2.2 The noncommutative relative distance zeta functional
where $\mu_{\tau}\in\mathcal{M}(P(H))$ is a Borel probability measure on $P(H):=\partial B_{1}^{H}(0)/\{\pm 1\}$; and we denote $\langle w,X\,w\rangle_{H}:=\langle r,X\,r\rangle_{\tilde{H}}$ whenever $r\in w$. When $n=\text{dim}_{\mathbb{R}}(H)$ can be possibly $\infty$, we may on the other hand consider the GNS representation of $\mathcal{L}(H)$ with $\tilde{H}:=\oplus_{l\in\mathcal{PS}(\mathcal{L}(H))}\,\mathcal{L}(H)/(l(a^{*}a)=0)$ (which is not separable), where we now denote $\mathcal{PS}(\mathcal{L}(H))$ the set of pure states of $\mathcal{L}(H)$. With that, we obtain the following representation of $\tau$ via the Choquet theorem:
@@
Hence, when $n=\text{dim}_{\mathbb{R}}(H)=2\,\text{dim}_{\mathbb{C}}(H)<\infty$, there is a canonical map between the coordinate vectors of P(H~)P($\tilde${H}) and the elements on $P(H)$, and the construction of $\tilde{H}$ is unnecessary (especially noticing that $\tilde{H}$ is way bigger). However, when $n=\text{dim}_{\mathbb{R}}(H)=\infty$, we cannot construct this map, as it is well-known that there are states $l$ such that $l(\mathcal{K}(H))=0$, where $\mathcal{K}(H)$ is the space of compact operators. (Meanwhile, as a possibly slightly confusing remark, if we start with an arbitrary $C^{*}$-algebra, we may take $H$ itself as the GNS representation of the $C^{*}$-algebra right from the start, and a state over that $C^{*}$-algebra can already be represented as an integral previously described (noting $n=\text{dim}_{\mathbb{R}}(H)=2\,\text{dim}_{\mathbb{C}}(H)=\infty$), i.e.,
$\tau(X)=\int_{P(H)}\langle w,X\,w\rangle_{H}\,\mu_{\tau}(dw)\,$,
@@
Motivated by the previous discussion as well as by the commutative case, let us define the following:
@@
We note that this definition is not restricted to the case when $n:=\text{dim}_{\mathbb{R}}(H)<\infty$, and the relative distance zeta functional can be explicitly written as
@@
Therefore our definition of (5) is not (interpreted in any sense as) a distance zeta function treating $\mathcal{L}(H)$ as a metric space. In fact, when $n$ is finite, our choice of having an infrimum over $\sigma(\mathcal{E}(X-Y))$ instead of a supremum compensates for the dimensionality effect within one variable $X_{i}$ from the aforementioned dimensionality argument. We postulate that we always have $\overline{\text{dim}}_{B}(\tau,\nu)$ on the strip $\{0\leq\Re(s)\leq d\}$, irrespective of whether $n$ is finite or infinite.
@@
Furthermore, if the relative box counting dimension of $\tau$ with respect to $\nu$, i.e. $\text{dim}_{B}(\tau,\nu)$, exists and $\text{dim}_{B}(\tau,\nu)<d$, and moreover, if $\mathcal{M}^{\text{dim}_{B}(\tau,\nu)}_{*}(\tau,\nu)>0$, then $\|\zeta_{\tau}(s,\nu)\|\rightarrow\infty$ when $s\rightarrow\text{dim}_{B}(\tau,\nu)^{+}$ for $s\in\mathbb{R}$.
@@
Given an arbitrary $\nu\in\tilde{\mathcal{M}}(\overline{\mathcal{A}\langle X\rangle})$, and for a fixed $\delta>0$, the $\delta$-tubular-neighborhood linear functional of $\tau\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ with respect to $\nu$ always satisfies $\text{spt}(\mu_{\nu^{\tau,\delta}})\subset\{(X,\xi)\in[\mathfrak{sa}\,\mathcal{L}(H)]^{d}\times P(\tilde{H}):d(X,\tau)\leq t\}$, and therefore $\zeta_{\tau}(s,\nu^{\tau,\delta})$ is well-defined and holomorphic whenever $\Re(s)>\overline{\text{dim}}_{B}(\mu,\nu^{\tau,\delta})$. Also, the above representation of $\zeta_{\tau}(s,\nu^{\tau,\delta})$ in the lemma shows that, given $g\in\overline{\mathcal{A}\langle X\rangle}$,
@@
We point out that it is now unclear if the relative complex dimension $\mathcal{D}_{\mathbb{C}}(\tau,\nu)$, as a union of $\mathcal{D}_{\mathbb{C}}(\tau,\nu)[g]$ over all $g\in\overline{\mathcal{A}\langle X\rangle}$, is in general a discrete set, unlike in the usual theory [42].
@@ 3.2 Decomposition rules
    Suppose $\tau\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ can be decomposed into $\tau=a_{1}\tau_{1}+a_{2}\tau_{2}$ with $\tau_{1},\tau_{2}\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ and $a_{1}+a_{2}=1$. Then, for $\nu\in\tilde{\mathcal{M}}(\overline{\mathcal{A}\langle X\rangle})$, then whenever $\overline{\text{dim}}_{B}(\tau_{2},\nu)<\overline{\text{dim}}_{B}(\tau,\nu)$, and for all $s\in\Omega_{\left[\zeta_{\tau_{1}}(\,\cdot\,,\nu)\right](g)}\bigcap\{s\in\mathbb{C}\,:\,\Re(s)>\overline{\text{dim}}_{B}(\tau_{2},\nu)\}$,
@@
    Suppose $\nu\in\tilde{\mathcal{M}}(\overline{\mathcal{A}\langle X\rangle})$ is given, and $\tau\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ can be decomposed into $\tau=a_{1}\tau_{1}+a_{2}\tau_{2}$ with $\tau_{1},\tau_{2}\in\tilde{\mathcal{P}}(\overline{\mathcal{A}\langle X\rangle})$ and $a_{1}+a_{2}=1$ such that for some $\epsilon>0$,
@@
where we know apriori that the right hand side of the second inequality is finite because $\left\|\nu\right\|<\infty$. Therefore, we have
@@ 3.3 Zeta functional over tensor states and decomposition rules
$\nu(f)=\int_{[\mathfrak{sa}\,\mathcal{L}(H)]^{\sum_{i=1}^{k}d_{i}}\times[P(\tilde{H})]^{k}}\,\left\langle f(X_{1},...,X_{k})\,\otimes_{i=1}^{k}\xi_{i}\,\,,\,\otimes_{i=1}^{k}\xi_{i}\right\rangle_{\tilde{H}^{\otimes^{k}}}\,\mu_{\nu}(d\xi_{1},...d\xi_{k},dX_{1},...,dX_{k})\,$,
@@
Furthermore, if the relative box counting dimension of $\tau$ with respect to $\nu$, i.e. $\text{dim}_{B}(\tau,\nu)$, exists and $\text{dim}_{B}(\tau,\nu)<d$, and moreover, if $\mathcal{M}^{\text{dim}_{B}(\tau,\nu)}_{*}(\tau,\nu)>0$, then $\|\zeta_{\tau}(s,\nu)\|\rightarrow\infty$ when $s\rightarrow\text{dim}_{B}(\tau,\nu)^{+}$ for $s\in\mathbb{R}$.
@@
$\qquad\qquad\qquad\qquad\qquad\prod_{i=1}^{k}\langle\xi_{i},g_{i}(X_{i})\,\xi_{i}\rangle_{\tilde{H}}\,(\otimes_{i=1}^{k}\mu_{\nu_{i}})(d\xi_{1},...d\xi_{k},dX_{1},...,dX_{k})\,.$
@@
The first part and third part of the lemma comes from an equivalence of norm in finite dimension